OpenStudy (anonymous):

An official volleyball measures 27 inches in circumference. find surface area.

OpenStudy (anonymous):

that's for the circumference.

OpenStudy (campbell_st):

so you need to find the radius so then the circumference formula is \[C = 2\pi r\] in your question you have \[27 = 2 \pi r\] make r the subject... when you have the value of r substitute it into the formula for surface area of a sphere \[SA = 4\pi r^2\]

OpenStudy (anonymous):

4pir^2 4*3.14(27/2)^2 12.56(13.5)^2 12.56*182.25 2,289.06

OpenStudy (anonymous):

thanks :)

OpenStudy (campbell_st):

I think your value for the radius is incorrect @heathernelly

OpenStudy (anonymous):

a bathroom floor tile measures 6 in on a side. if the whole floor contains 100 tiles, what is the area of the bathroom floor in square ft. Note: 1 ft = 12in

OpenStudy (campbell_st):

well go back to the circumference formula and make r the subject... correctly...

OpenStudy (anonymous):

is it? radius = 4.30?